# Yang-Mills Field in the $\kappa$-space-time

### v1

@@ 1 Introduction
As the existence of a fundamental length scale is not compatible with the special theory of relativity, principle of relativity needed to be modified. This lead to doubly special relativity (DSR) which admits a fundamental length scale and speed of light as universal constants [3]. The underlying space-time of DSR is known to be $\kappa$-space- time [4, 5, 6]. This $\kappa$-space-time also appears in the low energy limit of quantum gravity models [7]. The symmetries of $\kappa$-space-time is realised as Hopf algebra [4, 5]. Various issues in constructing quantum field theories in $\kappa$-space-time are addressed and resolved in [4, 5, 6, 7, 8, 9, 10, 11, 12, 16, 17, 18, 13, 14, 15, 19]
@@
$[\hat{x}^{i},\hat{x}^{j}]=0,~{}[\hat{x}^{0},\hat{x}^{i}]=ia\hat{x}^{i}$ (1.1)
where $a=\frac{1}{\kappa}$, by enlarging the scope of covariant generalisation of the Feynman’s approach to NC space-time. This requires the derivation of $\kappa$-deformed Wong’s equation. Using the $\kappa$-deformed Dirac Hamiltonian describing the interaction of isospin carrying fermion with non-abelian gauge field, we obtain the $\kappa$-deformed Wong’s equation. With the aid of this, we find the commutation relation between $\kappa$-deformed velocity and a Lie algebra valued functions in the $\kappa$-space-time. This also serves to provide the explicit form of gauge covariant derivative. Now replacing $[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]$ with a second rank, anti-symmetric tensor, in the Jacobi identity involving $\kappa$-deformed velocities, we obtain the equation satisfied by components of this second rank tensor. Employing Jacobi identities involving the deformed space-time coordinates and the generator of the Lie algebra, we find the second rank tensor introduced to be the Yang-Mills field strength in the $\kappa$-deformed space-time. Thus we derive the “homogeneous” Yang-Mills equation in the $\kappa$-deformed space-time. Using the $\kappa$-deformed field strength, we then write down the Lagrangian describing Yang-Mills field theory in the $\kappa$-deformed space-time. The remaining Yang-Mills equations follow as Euler-Lagrange equations from this $\kappa$-deformed Lagrangian. The force equation for a particle carrying isospin, interacting with Yang-Mills field is then derived.
@@
This paper is organised as follows: In sec.2, a brief sketch of Feynman’s approach of deriving Maxwell’s equation and Yang-Mills equations are given. We find Wong’s equation using the deformed Dirac Hamiltonian in $\kappa$-deformed space-time in sec.3. Using Wong’s equation, we derive the commutation relation between Lie algebra valued functions and deformed velocities. We also obtain an expression for modified gauge covariant derivative in the $\kappa$-deformed space-time. Using this we derive the Yang-Mills field equation in $\kappa$ deformed space-time in sec.4. We also obtain the explicit expression of $su(n)$ field strength in this section. We obtain the corresponding $\kappa$-deformed Lagrangian describing Yang-Mills theory and discuss its symmetry in sec.5. We derive an expression for the deformed Lorentz force experienced by a isospin carrying particle in the presence of $su(n)$ gauge field which is discussed in sec.6. We give the concluding remarks in sec.7.
@@ 3 Wong’s equation in $\kappa$-space-time
where $\gamma^{\mu}$ are the Dirac matrices satisfying the relations
$(\gamma^{0})^{2}=\mathbb{I};~{}~{}~{}\gamma^{0}\gamma^{i}=\alpha^{i};~{}~{}[\gamma^{\mu},\gamma^{\nu}]_{+}=2\eta^{\mu\nu}.$ (3.4)
@@
$p_{0}\rightarrow p_{0}-eA_{0},$
$p_{i}\rightarrow p_{i}-eA_{i},$ (3.8)
where $A_{\mu}$, $\mu=0,1,2,3$ are the components of the four-vector potential, we get $\kappa$-Dirac equation in presence of external non-abelian gauge field as,
$\gamma^{0}(p_{0}-eA_{0})\psi=-\bigg{[}\frac{a\gamma^{0}}{2\hbar}(p_{i}-eA_{i})^{2}+\gamma^{i}(p_{i}-eA_{i})-mc\bigg{]}\psi=0.$ (3.9)
Multiplying the above equation by $c\gamma^{0}$ and using the properties of gamma matrices, we re-express the above equation as
@@
$D_{0}\hat{x_{j}}$ $=$ $(\partial_{0}+\frac{ia}{2}\nabla^{2})(x_{j}-\frac{a}{\hbar}x_{j}p_{0})$ (3.15)
$=$ $\partial_{0}x_{j}(1-\frac{ap_{0}}{\hbar})-\frac{a}{\hbar c}x_{j}\frac{dp_{0}}{dt}$
$\cong$ $\partial_{0}x_{j}e^{-\frac{ap_{0}}{\hbar}}-\frac{a}{\hbar c}x_{j}\frac{dp_{0}}{dt},$
@@ 4 Yang-Mills equation in the $\kappa$-deformed space-time
In $\kappa$-deformed space-time, the position coordinates obey the commutation relation given by
Differentiating eq.(4.1) with respect to the parameter $\tau$ gives
@@
$[\hat{x}_{\mu},\dot{\hat{x}}_{\nu}]=-\frac{i\hbar}{m}\hat{\eta}_{\mu\nu},$ (4.3)
and choose $\eta_{\mu\nu}=(1,-1,-1,-1)$. Differentiating the above equation with respect to $\tau$ gives
$[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]+[\hat{x}_{\mu},\ddot{\hat{x}}_{\nu}]=0$ (4.4)
@@
$m[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]=-\frac{i\hbar e}{m}\hat{F}_{\mu\nu}$ (4.5)
Substituting this $\hat{F}_{\mu\nu}$ in the Jacobi identity, $[\hat{x}_{\mu},[\dot{\hat{x}}_{\nu},\dot{\hat{x}}_{\lambda}]]+[\dot{\hat{x}}_{\nu},[\dot{\hat{x}}_{\lambda},\hat{x}_{\mu}]]+[\dot{\hat{x}}_{\lambda},[\hat{x}_{\mu},\dot{\hat{x}}_{\nu}]]=0$ and using eq.(4.3), we find
$[\hat{x}_{\mu},\hat{F}_{\nu\lambda}]=0.$ (4.6)
Next consider the commutation relation
@@
Here as $[x_{0},I^{a}]=0$, the $a$-dependent term in Wong’s equation does not contribute to the above expression. Next we find the commutation relation between $\dot{\hat{x}}_{0}$ and Lie algebra valued functions of $\kappa$-deformed coordinates. For this we start with an arbitrary function $\phi^{a}(\hat{x}_{\mu})$ and consider its Taylor series expansion given by $\phi^{a}(\hat{x}_{\mu})=\phi^{a}(x_{\mu}-\frac{a}{\hbar}x_{\mu}p_{0})=\phi(x_{\mu})-\frac{a}{\hbar}x_{\nu}\partial^{\nu}\phi^{a}p_{0}$. Using this and the realisation of non-commutative coordinates in terms of commutative coordinates (see eq.(3.5)) and eq.(4.9), we get
@@
we find
$[\dot{\hat{x}}_{i},\phi^{a}(\hat{x}_{\mu})I^{a}]=(1-\frac{ap_{0}}{\hbar})\bigg{[}\frac{i\hbar}{m}(\delta^{ac}\partial_{i}-ef^{abc}A^{b}_{i}){\phi}^{c}I^{a}\bigg{]}$ (4.12)
Here all the $a$-dependent terms are coming from the realisation of $\hat{x}_{i}$ (see eq.(3.5)). As we are considering terms only upto first order in the deformation parameter $a$, $1-\frac{ap_{0}}{\hbar}$ can be taken as $e^{-\frac{ap_{0}}{\hbar}}$ and thus the above equation becomes
$[\dot{\hat{x}}_{i},\phi^{a}(\hat{x}_{\mu})I^{a}]=e^{-\frac{ap_{0}}{\hbar}}\frac{i\hbar}{m}\Big{(}\delta^{ac}\partial_{i}-ef^{abc}A^{b}_{i}\Big{)}{\phi}^{c}I^{a}=\frac{i\hbar}{m}\tilde{D}^{ac}_{i}\phi^{c}I^{a}$ (4.13)
where $\tilde{D}_{i}^{ac}\phi^{c}=e^{-\frac{ap_{0}}{\hbar}}(\delta^{ac}\partial_{i}-ef^{abc}A^{b}_{i}){\phi}^{c}$ is the spatial part of the gauge covariant derivative in $\kappa$-deformed space-time. Thus we obtain all the components of $\kappa$-deformed gauge covariant derivative. Using the Jacobi identity
$[\dot{\hat{x}}_{\mu},[\dot{\hat{x}}_{\nu},\dot{\hat{x}}_{\rho}]]+[\dot{\hat{x}}_{\nu},[\dot{\hat{x}}_{\rho},\dot{\hat{x}}_{\mu}]]+[\dot{\hat{x}}_{\rho},[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]]=0$ (4.14)
and substituting
$[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]=-\frac{i\hbar e}{m^{2}}\hat{F}_{\mu\nu}$
in the above, we find that the $\hat{F}_{\mu\nu}$ satisfies
@@
It remains to be shown that $\hat{F}_{ij}$ and $\hat{F}_{0i}$ are indeed $\kappa$-deformed Yang-Mills field strength. To show this, we start from the commutation relation between $\dot{\hat{x}}_{0}$ and $I^{a}$ given in eq.(4.9). Taking commutator of eq.(4.9) with $\dot{\hat{x}}_{i}$ gives,
@@
Using the Jacobi idenity for the first term and eq.(3.5), the above equation becomes
00 0 $=$ $[\dot{\hat{x}}_{0},[\dot{\hat{x}}_{i},I^{a}]]+[I^{a},[\dot{\hat{x}}_{0},\dot{\hat{x}}_{i}]]-\frac{i\hbar e}{m}f^{abc}\Big{(}[\dot{x}_{i}-\frac{a}{\hbar}\dot{x}_{i}p_{0},A^{b}_{0}I^{c}]\Big{)}$ (4.18)
$=$ $[\dot{x}_{0},[\dot{\hat{x}}_{i},I^{a}]]-\frac{i\hbar e}{m^{2}}[I^{a},\hat{F}^{b}_{0i}I^{b}]-\frac{i\hbar e}{m}f^{abc}\Big{(}[\dot{x}_{i}-\frac{a}{\hbar}\dot{x}_{i}p_{0},A^{b}_{0}I^{c}]\Big{)}$
In writing the second term in the above equation, we have used the definition of $\hat{F^{a}_{0i}}$ given in eq.(4.5). Using eq.(4.11) and eq.(3.1), we get the expression for $\hat{F^{a}_{0i}}$ as
@@
Here we have taken $1-\frac{ap_{0}}{\hbar}$ to be $e^{-\frac{ap_{0}}{\hbar}}$, as we are considering corrections only upto first order term in $a$. Here $F^{a}_{0i}=\partial_{0}A^{a}_{i}-\partial_{i}A^{a}_{0}-ef^{abc}A^{b}_{0}A^{c}_{i}$, is the commutative “Electric field”.
@@
By applying the Jacobi identity in the first term,
$0=$ $~{}[\dot{\hat{x_{i}}},[\dot{\hat{x_{j}}},I^{a}]]+[I^{a},[\dot{\hat{x_{i}}},\dot{\hat{x_{j}}}]]-\frac{i\hbar e}{m}f^{abc}([\dot{\hat{x_{j}}},A^{b}_{i}]I^{c}+A^{b}_{i}[\dot{\hat{x_{j}}},I^{c}])+\frac{iae}{m}f^{abc}([\dot{\hat{x_{j}}},A^{b}_{i}I^{c}p_{0}])$
$=$ $~{}[\dot{\hat{x_{i}}},[\dot{\hat{x_{j}}},I^{a}]]-\frac{i\hbar e}{m^{2}}\hat{F}^{b}_{ij}[I^{a},I^{b}]-\frac{i\hbar e}{m}f^{abc}([\dot{\hat{x_{j}}},A^{b}_{i}]I^{c}+A^{b}_{i}[\dot{\hat{x_{j}}},I^{c}])+\frac{iae}{m}f^{abc}([\dot{\hat{x_{j}}},A^{b}_{i}I^{c}p_{0}])$ (4.21)
In writing the second term above, we have used eq.(4.5). Next using eq.(4.11) and eq.(3.1) in the right hand side of the above equation and simplifying, we get
@@
where non-commutative correction is approximated as the exponential and $F^{a}_{ij}=\partial_{i}A^{a}_{j}-\partial_{j}A^{a}_{i}-ef^{abc}A^{b}_{i}A^{c}_{j}$ is the commutative “magnetic” field.
@@
$([\tilde{D}_{0},\tilde{D}_{i}]\phi)^{a}$ $=$ $-ef^{abc}F^{b}_{0i}\phi^{c}(1-\frac{ap_{0}}{\hbar})$ (4.23)
$=$ $-ef^{abc}F^{b}_{0i}\phi^{c}e^{-{\frac{ap_{0}}{\hbar}}}$
@@
$([\tilde{D}_{i},\tilde{D}_{j}]\phi)^{a}$ $=$ $-ef^{abc}F^{b}_{ij}\phi^{c}(1-\frac{2ap_{0}}{\hbar})$ (4.24)
$=$ $-ef^{abc}F^{b}_{ij}\phi^{c}e^{-{\frac{2ap_{0}}{\hbar}}}$

### v4

@@ 1 Introduction
As the existence of a fundamental length scale is not compatible with the special theory of relativity, principle of relativity needed to be modified. This lead to doubly special relativity (DSR) which admits a fundamental length scale and speed of light as universal constants [3]. The underlying space-time of DSR is known to be $\kappa$-space- time [4, 5, 6]. This $\kappa$-space-time also appears in the low energy limit of quantum gravity models [7]. The symmetries of $\kappa$-space-time is realised as Hopf algebra [4, 5]. Various issues in constructing quantum field theories in $\kappa$-space-time are addressed and resolved in [4, 5, 6, 7, 8, 9, 10, 11, 12, 16, 17, 18, 13, 14, 15, 19, 20, 21]
@@
with $a_{0}=a=\frac{1}{\kappa}$, $a_{i}=0$, by enlarging the scope of covariant generalisation of the Feynman’s approach to NC space-time. This requires the derivation of $\kappa$-deformed Wong’s equation. Using the $\kappa$-deformed Dirac Hamiltonian describing the interaction of isospin carrying fermion with non-abelian gauge field, we obtain the $\kappa$-deformed Wong’s equation. With the aid of this, we find the commutation relation between $\kappa$-deformed velocity and a Lie algebra valued functions in the $\kappa$-space-time. This also serves to provide the explicit form of gauge covariant derivative. Now replacing $[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]$ with a second rank, anti-symmetric tensor, in the Jacobi identity involving $\kappa$-deformed velocities, we obtain the equation satisfied by components of this second rank tensor. Employing Jacobi identities involving the deformed space-time coordinates and the generator of the Lie algebra, we find the second rank tensor introduced to be the Yang-Mills field strength in the $\kappa$-deformed space-time. Thus we derive the “homogeneous” Yang-Mills equation in the $\kappa$-deformed space-time. Using the $\kappa$-deformed field strength, we then write down the Lagrangian describing Yang-Mills field theory in the $\kappa$-deformed space-time. The remaining Yang-Mills equations follow as Euler-Lagrange equations from this $\kappa$-deformed Lagrangian. The force equation for a particle carrying isospin, interacting with Yang-Mills field is then derived.
@@
After quickly recalling the derivations of Maxwell’s equations and Yang-Mills equation using Feynman’s approach, we present the derivation of $\kappa$-deformed Wong’s equation. We then derive Yang-Mills equation in the $\kappa$-deformed space-time. Commutator between the $\kappa$-deformed velocity and Lie algebra valued functions of $\kappa$-deformed space-time is an integral part of this derivation and is calculated using $\kappa$-deformed Wong’s equation. Using the explicit form of the Yang-Mills field strength in the $\kappa$-deformed space-time obtained, we then construct the Lagrangian describing $\kappa$-deformed Yang-Mills theory. We show that this model is invariant under SU(N) transformations. We then derive the force experienced by an isospin carrying particle moving in presence of Yang-Mills fields.
@@ 3 Wong’s equation in $\kappa$-space-time
where $\gamma^{\mu}$ are the Dirac matrices satisfying the relations $(\gamma^{0})^{2}=\mathbb{I};~{}~{}~{}\gamma^{0}\gamma^{i}=\alpha^{i};~{}~{}[\gamma^{\mu},\gamma^{\nu}]_{+}=2\eta^{\mu\nu}$.
@@
$p_{0}\rightarrow p_{0}-eA_{0},~{}~{}p_{i}\rightarrow p_{i}-eA_{i},$ (3.7)
where $A_{\mu}$, $\mu=0,1,2,3$ are the components of the four-vector potential, and simplifying we get $\kappa$-deformed Dirac equation in presence of external non-abelian gauge field as,
@@
$D_{0}\hat{x}_{j}=(\partial_{0}+\frac{ia}{2}\nabla^{2})(x_{j}-\frac{a}{\hbar}x_{j}p_{0}),~{}~{}=\partial_{0}x_{j}(1-\frac{ap_{0}}{\hbar})-\frac{a}{\hbar c}x_{j}\frac{dp_{0}}{dt}~{}~{}\cong\partial_{0}x_{j}e^{-\frac{ap_{0}}{\hbar}}-\frac{a}{\hbar c}x_{j}\frac{dp_{0}}{dt},$ (3.13)
@@ 4 Yang-Mills equation in the $\kappa$-deformed space-time
In $\kappa$-deformed space-time, the position coordinates obey the commutation relation given by eq.(1.1). Differentiating eq.(1.1) with respect to the parameter $\tau$ gives
@@
$[\hat{x}_{\mu},\dot{\hat{x}}_{\nu}]=-\frac{i\hbar}{m}\hat{\eta}_{\mu\nu}.$ (4.2)
Differentiating the above equation with respect to $\tau$ gives
$[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]+[\hat{x}_{\mu},\ddot{\hat{x}}_{\nu}]=-\frac{i\hbar}{m}\frac{d\hat{\eta}_{\mu\nu}}{d\tau}$ (4.3)
@@
$m[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]=-\frac{i\hbar e}{m}\hat{F}_{\mu\nu}+\hat{S}_{\mu\nu}$ (4.4)
where $\hat{S}_{\mu\nu}=-\frac{i\hbar}{m}\frac{d\hat{\eta}_{\mu\nu}}{d\tau}=-\frac{i\hbar}{m}\frac{d\delta\hat{\eta}_{\mu\nu}}{d\tau}$ as the commutative metric is constant in $\tau$. Thus $\hat{S}_{\mu\nu}$ has no commutative part. Next consider the commutation relation
@@
Here as $[x_{0},I^{a}]=0$, the $a$-dependent term in Wong’s equation does not contribute to the above expression. Next we find the commutation relation between $\dot{\hat{x}}_{0}$ and Lie algebra valued functions of $\kappa$-deformed coordinates. For this we start with an arbitrary function $\phi^{a}(\hat{x}_{\mu})$ and consider its Taylor series expansion given by $\phi^{a}(\hat{x}_{\mu})=\phi^{a}(x_{\mu}-\frac{a}{\hbar}x_{\mu}p_{0})=\phi(x_{\mu})-\frac{a}{\hbar}x_{\nu}\partial^{\nu}\phi^{a}p_{0}$. Using this and the realisation of non-commutative coordinates in terms of commutative coordinates (see eq.(3.4) and eq.(4.7)), we get
@@
${\rm~{}we~{}find~{}~{}}[\dot{\hat{x}}_{i},\phi^{a}(\hat{x}_{\mu})I^{a}]=(1-\frac{ap_{0}}{\hbar})\frac{i\hbar}{m}\Big{(}\delta^{ac}\partial_{i}-ef^{abc}A^{b}_{i}\Big{)}{\phi}^{c}I^{a}=\frac{i\hbar}{m}\tilde{D}^{ac}_{i}\phi^{c}I^{a}$ (4.10)
where $\tilde{D}_{i}^{ac}\phi^{c}=e^{-\frac{ap_{0}}{\hbar}}(\delta^{ac}\partial_{i}-ef^{abc}A^{b}_{i}){\phi}^{c}$ is the spatial part of the gauge covariant derivative in $\kappa$-deformed space-time. Here all the $a$-dependent terms are coming from the realisation of $\hat{x}_{i}$ (see eq.(3.4)). As we are considering terms only upto first order in the deformation parameter $a$, $1-\frac{ap_{0}}{\hbar}$ is taken as $e^{-\frac{ap_{0}}{\hbar}}$. Thus we obtain all the components of $\kappa$-deformed gauge covariant derivative. Using the Jacobi identity $[\dot{\hat{x}}_{\mu},[\dot{\hat{x}}_{\nu},\dot{\hat{x}}_{\rho}]]+[\dot{\hat{x}}_{\nu},[\dot{\hat{x}}_{\rho},\dot{\hat{x}}_{\mu}]]+[\dot{\hat{x}}_{\rho},[\dot{\hat{x}}_{\mu},\dot{\hat{x}}_{\nu}]]=0$ and substituting eq.(4.4) in the above, we find that the $\hat{F}_{\mu\nu}$ satisfies
@@
where we used $[\dot{\hat{x}}^{\mu},\frac{d\delta\hat{\eta}^{\nu\rho}}{d\tau}]=0$ as the $\kappa$-deformed corrections to the metric is known to be independent of $\hat{x}^{\mu}$ and depend only on $p^{\mu}$ [34, 35, 36]. It remains to be shown that $\hat{F}_{ij}$ and $\hat{F}_{0i}$ are indeed $\kappa$-deformed Yang-Mills field strength. To show this, we start from the commutation relation between $\dot{\hat{x}}_{0}$ and $I^{a}$ given in eq.(4.7). Taking commutator of eq.(4.7) with $\dot{\hat{x}}_{i}$ gives,
@@
Using the Jacobi idenity to re-express the first term as $[\dot{\hat{x}}_{0},[\dot{\hat{x}}_{i},I^{a}]]+[I^{a},[\dot{\hat{x}}_{0},\dot{\hat{x}}_{i}]]$ and eq.(3.4), the above equation becomes
00 0 $=$ $[\dot{x}_{0},[\dot{\hat{x}}_{i},I^{a}]]-\frac{i\hbar e}{m^{2}}[I^{a},\hat{F}^{b}_{0i}I^{b}]-\frac{i\hbar e}{m}f^{abc}\Big{(}[\dot{x}_{i}-\frac{a}{\hbar}\dot{x}_{i}p_{0},A^{b}_{0}I^{c}]\Big{)}$ (4.14)
In writing the second term in the above equation, we have used the eq.(4.4). Using eq.(4.9) and eq.(3.1), we get the expression for $\hat{F}^{a}_{0i}$ as
@@
As earlier we have taken $1-\frac{ap_{0}}{\hbar}$ to be $e^{-\frac{ap_{0}}{\hbar}}$. Here $F^{a}_{0i}=\partial_{0}A^{a}_{i}-\partial_{i}A^{a}_{0}-ef^{abc}A^{b}_{0}A^{c}_{i}$, is the commutative “Electric field”.
@@
As earlier applying the Jacobi identity in the first term and using eq.(4.4), we find
$0=$ $~{}[\dot{\hat{x}}_{i},[\dot{\hat{x}}_{j},I^{a}]]-\frac{i\hbar e}{m^{2}}\hat{F}^{b}_{ij}[I^{a},I^{b}]-\frac{i\hbar e}{m}f^{abc}([\dot{\hat{x}}_{j},A^{b}_{i}]I^{c}+A^{b}_{i}[\dot{\hat{x}}_{j},I^{c}])+\frac{iae}{m}f^{abc}([\dot{\hat{x}}_{j},A^{b}_{i}I^{c}p_{0}])$ (4.17)
Next using eq.(4.9) and eq.(3.1) in the right hand side of the above equation and simplifying, we get
@@
where NC correction is approximated as the exponential and $F^{a}_{ij}=\partial_{i}A^{a}_{j}-\partial_{j}A^{a}_{i}-ef^{abc}A^{b}_{i}A^{c}_{j}$ is the commutative “magnetic” field.
@@
$([\tilde{D}_{0},\tilde{D}_{i}]\phi)^{a}=-ef^{abc}F^{b}_{0i}\phi^{c}(1-\frac{ap_{0}}{\hbar})=-ef^{abc}F^{b}_{0i}\phi^{c}e^{-{\frac{ap_{0}}{\hbar}}}$ (4.19)
@@
$([\tilde{D}_{i},\tilde{D}_{j}]\phi)^{a}=-ef^{abc}F^{b}_{ij}\phi^{c}(1-\frac{2ap_{0}}{\hbar})=-ef^{abc}F^{b}_{ij}\phi^{c}e^{-{\frac{2ap_{0}}{\hbar}}}$ (4.20)